# MRIR: Integrating Multimodal Insights for Diffusion-based Realistic Image Restoration

### v1

@@ IV-B Experimental Results
Qualitative Comparison. We first provide the qualitative comparison in Fig 3 and Fig 4. For the synthetic datasets, as shown in Fig 3, GAN-based methods often produce smoother restoration results. When the image degradation is severe, they may fail to generate accurate results. In contrast, diffusion-based models can generate more detailed results. However, since StableSR and DiffBIR do not utilize semantic information to guide the diffusion model, their restoration results often lack richness in details and exhibit unclear textures. For example, the face structure (row 3) is unclear and the texture of the pineapple (row 4) is blurry. Although PASD and SeeSR take semantic information into consideration, they also produce incorrect restoration results due to inaccurate semantic predictions. As shown in Fig 3, PASD mistakenly identifies the pineapple as another fruit, resulting in an erroneous generated output (row 4) while SeeSR fails to correctly identify the building, resulting in the generation of rocks in place of the building (row 2). Additionally, benefiting from the image understanding capability of MLLM, our method is able to predict accurate semantic information from low-resolution images. Specifically, as shown in row 1, the restoration results of other methods do not include dewdrops on the flowers, while only our method’s restoration results contain dewdrops. Moreover, the structure and texture of the flowers generated by our method are more defined and clearer. To be honest, due to significant information loss in degraded images, achieving image restoration results that are comparable to ground truth is challenging and the restored images obtained through our method represent high-quality and plausible solutions.
@@
For the real-world datasets, as shown in Fig 4, our method also generates results with reasonable structure, accurate semantics, and rich details such as the text on the chip (row 1), the string structure of the violin (row 2), the eye details (row 3), the texture details of Spider-Man’s facial details (row 4) and the structure and clarity of the succulent (row 5).
@@
In conclusion, our method demonstrates superior performances in terms of both the rationality of structures and the clarity of details, thanks to the powerful image understanding and reasoning capabilities of MLLM, as well as the application of pixel control.
Quantitative Comparison. We then provide the quantitative comparison in Table I and Table II. Table I shows the quantitative results on the synthetic datasets and Table II shows the quantitative results on the real-world datasets. Compared with diffusion-based methods, our method achieves the best results in MANIQA, MUSIQ, and CLIPIQA on all five datasets, except for the score of MUSIQ on RealLR200. What’s more, our method outperforms other state-of-the-art methods in LPIPS, FID, and DISTS on the DIV2K-Val dataset and LSDIR-Val dataset. Compared with GAN-based methods, diffusion-based methods tend to get lower PSNR and SSIM scores. This is mainly because diffusion-based methods tend to generate more realistic details but at the expense of fidelity. While PSNR and SSIM are commonly used objective metrics for image quality assessment, they have limitations in capturing perceptual quality and fine details, as mentioned in [51]. Overall, our method achieves higher scores in no-reference metrics while maintaining competitive performance in full-reference measures.
User Study. We conduct a user study with 20 participants to assess various methods through subjective evaluation. We compared our method with other six competitive restoration methods (BSRGAN, RealESRGAN, StableSR, DiffBIR, PASD, and SeeSR). We provided LQ images as references and asked the participants to select the best restoration result from the seven methods mentioned above. We randomly selected 26 images from synthetic and real-world datasets for testing and received evaluations from 20 participants, resulting in a total of 520 votes (26 images × 20 participants). As shown in Fig 5, our method significantly outperformed all other six competing methods, receiving 52.69% of the votes, surpassing the second-place method with 38.84% support.
@@
Figure 6: Visual comparison for the ablation study of different textual level settings. (1) is the result with null prompt, (2) is the result with BLIP prompt, (3) is the result with null image embedding, (4) is the result of replacing the refine layer with only one linear layer, (5) is the result with null prompt and null image embedding, and (6) is the result of our method.
TABLE III: Ablation studies of different prompts on DIV2K-Val and RealSR datasets.
@@
Experiments PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
(1) 19.56 41.40 0.3806 66.20 0.6606 23.66 136.77 0.2936 68.82 0.6637
(2) 19.38 38.75 0.3864 66.11 0.6582 23.57 131.62 0.2905 68.10 0.6451
(3) 19.09 39.28 0.4001 65.77 0.6646 23.31 134.41 0.2970 69.53 0.6591
(4) 19.28 39.32 0.3903 66.35 0.6723 23.69 126.54 0.2827 66.87 0.6224
(5) 19.50 45.63 0.3973 64.90 0.6477 23.61 143.57 0.2993 69.53 0.6723
Ours 19.18 38.07 0.3840 69.27 0.7088 23.76 123.21 0.2865 69.81 0.6765
$L_{RGB}$ $L_{FFT}$ DIV2K-Val RealSR
PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
19.19 38.80 0.3954 67.03 0.6765 23.07 133.60 0.3025 68.81 0.6564
19.16 39.24 0.3846 68.63 0.7087 23.70 125.80 0.2995 69.39 0.6644
19.15 39.14 0.3893 68.19 0.7060 23.03 123.34 0.3028 70.21 0.6817
19.18 38.07 0.3840 69.27 0.7091 23.76 123.21 0.2865 69.81 0.6765
Image Pixel DIV2K-Val RealSR
Attention Attention PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
18.87 44.41 0.4223 68.22 0.7085 22.57 139.41 0.3255 70.42 0.6740
19.29 37.80 0.3897 65.15 0.6539 23.62 124.98 0.2787 67.26 0.6093
19.63 45.45 0.3909 65.73 0.6416 23.81 137.25 0.2898 68.47 0.6393
19.18 38.07 0.3840 69.27 0.7091 23.76 123.21 0.2865 69.81 0.6765
@@ IV-C Ablation Study
To validate the effectiveness of the proposed method in detail, we conducted ablation experiments from the perspectives of textual level control, pixel level control, and attention mechanism.
Firstly, we analyze the influence of textual level control, specifically the impact of text embedding and image embedding. We conduct experiments considering the following five cases:
@@
The results are shown in Table III and Fig 6. Table III showcases the significant enhancement in metrics such as FID, LPIPS, MUSIQ, and CLLIPIQA by incorporating text embedding and image embedding. What’s more, the PSNR metric of our method remains at a comparable level. The quantitative metrics indicate that our approach excels in both semantic and texture generation, resulting in more realistic and accurate outputs. Additionally, it can be observed in Fig 6 that setting both the text embedding and image embedding to null leads to the loss of texture details. Additionally, using BLIP-generated prompts as a replacement for MLLM-generated prompts may result in suboptimal restoration outcomes due to potential inaccuracies in the prompts generated from LQ images, such as the incorrectly generated airplane in Fig 1 (b) and the engraved pattern on the stone in Fig 6 (2). Furthermore, when replacing the refine layer with only one linear layer, the lack of correction processes significantly decreases the restoration performance, specifically regarding the clarity of the statue texture highlighted by the red box in Fig 6, and leads to a decline in evaluation metrics as well.
@@
Figure 8: Visual comparison for the ablation study of different attention settings. With the image attention and pixel attention, the tire tread pattern is more realistic and clearer.
@@
Next, we evaluate the impact of pixel level loss constraints on the restoration results. The results are presented in Table IV and Fig 7. It can be observed that removing $L_{RGB}$ and $L_{FFT}$ leads to an overall decline in the generation quality. Although removing $L_{RGB}$ may yield gains in terms of MUSIQ and CLIPIQA metrics on the RealSR dataset, the generated details may contain artifacts and the structural details may not be sufficiently clear. As shown in Fig 7, removing $L_{RGB}$ and $L_{FFT}$ results in artifacts on the person’s hair, and the details of the clothing appear blurry. Removing $L_{FFT}$ leads to an unclear structure of the belt on the person’s clothing. Removing $L_{RGB}$ causes overall blurry artifacts on the person’s clothing. In contrast, our method performs better in terms of preserving the shape contours of the person and capturing clothing details.
@@
Lastly, we evaluate the effectiveness of the introduced image cross attention and pixel attention. The experimental results are shown in Table V and Fig 8. In terms of quantitative metrics, it is evident that the removal of image attention and pixel attention leads to a significant decrease in PSNR, indicating a decline in the accuracy of the generated structural details. Simultaneously, the individual removal of image attention and pixel attention impacts metrics such as FID and MUSIQ, which assess the quality of the generated images. Our method achieves a trade-off between realism and fidelity. Visually, it can be observed in Fig 8 that when removing image attention, the texture of the tire is not very clear. When removing pixel attention, although the texture of the tire is clear, the structural details of the tire’s texture are incomplete. When removing image attention and pixel attention, both the texture structure and clarity are compromised. Our method achieves better results in terms of structure and clarity thanks to both image attention and pixel attention being employed.
V Conclusion
@@
In this paper, we propose a novel multimodal-perception approach for realistic image restoration based on the diffusion model. We consider both textual level and pixel level perspectives to solve the restoration problem. Regarding textual level information, we utilize the multimodal large language model to infer reliable semantic text representations from low-quality images. Additionally, we design the Refine Layer to generate the corrected image embeddings, which collaborate with MLLM to achieve a balance between fidelity and realism. For structural control in generating images, we employ the Pixel-level Processor and ControlNet for pixel control. Lastly, we integrate the control information and the diffusion model using the attention mechanism, enabling high-fidelity and high-realism image restoration. Our method demonstrates state-of-the-art performance on multiple datasets both qualitatively and quantitatively.

### v2

@@ IV-B Experimental Results
Qualitative Comparison. We first provide the qualitative comparison in Fig 4 and Fig 5. For the synthetic datasets, as shown in Fig 4, GAN-based methods often produce smoother restoration results. When the image degradation is severe, they may fail to generate accurate results. In contrast, diffusion-based models can generate more detailed results. However, since StableSR and DiffBIR do not utilize semantic information to guide the diffusion model, their restoration results often lack richness in details and exhibit unclear textures. For example, the face structure (row 3) is unclear and the texture of the pineapple (row 4) is blurry. Although PASD and SeeSR take semantic information into consideration, they also produce incorrect restoration results due to inaccurate semantic predictions. As shown in Fig 4, PASD mistakenly identifies the pineapple as another fruit, resulting in an erroneous generated output (row 4) while SeeSR fails to correctly identify the building, resulting in the generation of rocks in place of the building (row 2). Additionally, benefiting from the image understanding capability of MLLM, our method is able to predict accurate semantic information from low-resolution images. Specifically, as shown in row 1, the restoration results of other methods do not include dewdrops on the flowers, while only our method’s restoration results contain dewdrops. Moreover, the structure and texture of the flowers generated by our method are more defined and clearer. To be honest, due to significant information loss in degraded images, achieving image restoration results that are exactly the same as ground truth is challenging. The restored images obtained through our method only represent high-quality and plausible solutions.
@@
For the real-world datasets, as shown in Fig 5, our method also generates results with reasonable structure, accurate semantics, and rich details such as the text on the chip (row 1), the string structure of the violin (row 2), the eye details (row 3), the texture details of Spider-Man’s facial details (row 4) and the structure and clarity of the succulent (row 5). In contrast, SUPIR often generates excessive details, such as wrinkles in the eyes (row 3), surface of the succulent (row 5), which are not faithful to the input images.
@@
In conclusion, our method demonstrates superior performances in terms of both the rationality of structures and the clarity of details, thanks to the powerful image understanding and reasoning capabilities of MLLM, as well as the application of image embedding and structure control.
@@
TABLE III: Ablation studies of different semantic prompts on DIV2K-Val and RealSR datasets.
@@
Experiments PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
(1) 19.56 41.40 0.3806 0.8814 66.20 0.6606 23.66 136.77 0.2936 0.8854 68.82 0.6637
(2) 19.38 38.75 0.3864 0.8898 66.11 0.6582 23.57 131.62 0.2905 0.8881 68.10 0.6451
(3) 19.09 39.28 0.4001 0.8915 65.77 0.6646 23.31 134.41 0.2970 0.8880 69.53 0.6591
(4) 19.28 39.32 0.3903 0.8955 66.35 0.6723 23.69 126.54 0.2827 0.8939 66.87 0.6224
(5) 19.50 45.63 0.3973 0.8663 64.90 0.6477 23.61 143.57 0.2993 0.8817 69.53 0.6723
Ours 19.18 38.07 0.3840 0.8982 69.27 0.7088 23.76 123.21 0.2865 0.8956 69.81 0.6765
Quantitative Comparison. We then provide the quantitative comparison in Table I and Table II. Table I shows the quantitative results on the synthetic datasets and Table II shows the quantitative results on the real-world datasets. Compared with diffusion-based methods, our method achieves the best results in CLIP-I, MANIQA, MUSIQ, and CLIPIQA on all five datasets, except for the score of MUSIQ on RealLR200. This demonstrates the effectiveness of our proposed semantic prior extraction module and the superiority of our approach in terms of semantic accuracy and perceptual quality. What’s more, our method outperforms other state-of-the-art methods in LPIPS, FID, and DISTS on the DIV2K-Val dataset and LSDIR-Val dataset. Compared with GAN-based methods, diffusion-based methods tend to getlower PSNR and SSIM scores. This is mainly because diffusion-based methods tend to generate more realistic details but at the expense of fidelity. While PSNR and SSIM are commonly used objective metrics for image quality assessment, they have limitations in capturing perceptual quality and fine details, as mentioned in [42]. However, for the structural aspect, our method also obtains a relatively high SSIM score on all datasets compared with other diffusion-based methods, demonstrating the structure-preserving ability of the proposed method with the structure prior. Overall, our method achieves higher scores in no-reference metrics while maintaining competitive performance in reference metrics.
User Study. We conduct a user study with 20 participants to assess various methods through subjective evaluation. We compared our method with other six competitive restoration methods (BSRGAN, RealESRGAN, StableSR, DiffBIR, PASD, and SeeSR). We provided LQ images as references and asked the participants to select the best restoration result from the seven methods mentioned above. We randomly selected 26 images from synthetic and real-world datasets for testing and received evaluations from 20 participants, resulting in a total of 520 votes (26 images × 20 participants). As shown in Fig 6, our method significantly outperformed all other six competing methods, receiving 52.69% of the votes, 38.84% higher than the second place.
Figure 7: Visual comparison for the ablation study of different semantic prompts settings. (1) is the result with null prompt, (2) is the result with BLIP prompt, (3) is the result with null image embedding, (4) is the result of replacing the Refine Layer with only one linear layer, (5) is the result with null prompt and null image embedding, and (6) is the result of our method.
@@ IV-C Ablation Study
To validate the effectiveness of the proposed method in detail, we conducted ablation experiments from the perspectives of semantic prior, structure prior, and attention module. Additionly, we also validated the directional restoration capability of our proposed method driven by textual prompts.
Firstly, we analyze the influence of semantic prior, specifically the impact of text embedding and image embedding. We conduct experiments considering the following five cases:
@@
$\mathcal{L}_{RGB}$ $\mathcal{L}_{FFT}$ DIV2K-Val RealSR
PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
19.19 38.80 0.3954 0.8946 67.03 0.6765 23.07 133.60 0.3025 0.8851 68.81 0.6564
19.16 39.24 0.3846 0.8964 68.63 0.7087 23.70 125.80 0.2995 0.8902 69.39 0.6644
19.15 39.14 0.3893 0.8935 68.19 0.7060 23.03 123.34 0.3028 0.8882 70.21 0.6817
19.18 38.07 0.3840 0.8982 69.27 0.7091 23.76 123.21 0.2865 0.8956 69.81 0.6765
Image Strcuture DIV2K-Val RealSR
Attention Attention PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
18.87 44.41 0.4223 0.8803 68.22 0.7085 22.57 139.41 0.3255 0.8736 70.42 0.6740
19.29 37.80 0.3897 0.8868 65.15 0.6539 23.62 124.98 0.2787 0.8918 67.26 0.6093
19.63 45.45 0.3909 0.8678 65.73 0.6416 23.81 137.25 0.2898 0.8789 68.47 0.6393
19.18 38.07 0.3840 0.8982 69.27 0.7091 23.76 123.21 0.2865 0.8956 69.81 0.6765
@@
Figure 9: Visual comparison for the ablation study of different attention settings. With the image attention and structure attention, the tire tread pattern is more realistic and clearer.
Figure 10: Visual comparison for directional restoration capability of different prompt. Users can modify the details of the restoration results as needed.
The results are shown in Table III and Fig 7. Table III showcases the significant enhancement in metrics such as FID, LPIPS, CLIP-I, MUSIQ, and CLIPIQA by incorporating text embedding and image embedding. What’s more, the PSNR metric of our method remains at a comparable level. Especially when setting both the image embedding and text embedding to null will lead to a significant decrease in CLIP-I, which fully proves the effectiveness of our proposed semantic prior. The quantitative metrics indicate that our approach excels in both semantic and texture generation, resulting in more realistic and accurate outputs. Additionally, it can be observed in Fig 7 that setting both the text embedding and image embedding to null leads to the loss of texture details. Additionally, using BLIP-generated prompts as a replacement for MLLM-generated prompts may result in suboptimal restoration outcomes due to potential inaccuracies in the prompts generated from LQ images, such as the incorrectly generated airplane in Fig 3 (b) and the engraved pattern on the stone in Fig 7 (2). Furthermore, when replacing the Refine Layer with only one linear layer, the lack of correction processes significantly decreases the restoration performance, specifically regarding the clarity of the statue texture highlighted by the red box in Fig 7, and leads to a decline in evaluation metrics as well.
Figure 11: More examples to show the generalizability of the proposed method. Our method achieves great performance in various scenes.
TABLE VI: Analysis of the model’s complexity. All methods are tested with an input image of size 512 × 512, and the inference time is measured on an A100 GPU.
Methods Real-ESRGAN [17] StableSR [31] DiffBIR [32] PASD [33] SeeSR [34] SUPIR [42] Ours
One-stage Training
Trainable Params (M) 16.70 153.27 378.94 609.52 751.66 - 997.80
Inference Time (s) 0.08 14.50 4.10 3.38 4.30 17.68 8.82
@@
Next, we evaluate the impact of loss constraints for stucture prior extraction on the restoration results. The results are presented in Table IV and Fig 8. It can be observed that removing $\mathcal{L}_{RGB}$ and $\mathcal{L}_{FFT}$ leads to an overall decline in the generation quality. Although removing $\mathcal{L}_{RGB}$ may yield gains in terms of MUSIQ and CLIPIQA metrics on the RealSR dataset, the generated details may contain artifacts and the structural details may not be sufficiently clear. As shown in Fig 8, removing $\mathcal{L}_{RGB}$ and $\mathcal{L}_{FFT}$ results in artifacts on the person’s hair, and the details of the clothing appear blurry. Removing $\mathcal{L}_{FFT}$ leads to an unclear structure of the belt on the person’s clothing. Removing $\mathcal{L}_{RGB}$ causes overall blurry artifacts on the person’s clothing. In contrast, our method performs better in terms of preserving the shape contours of the person and capturing clothing details.
@@
Then, we evaluate the effectiveness of the introduced image cross attention and structure attention in Prior-Guided Attention Module. The experimental results are shown in Table V and Fig 9. In terms of quantitative metrics, it is evident that the removal of image attention and structure attention leads to a significant decrease in PSNR, indicating a decline in the accuracy of the generated structural details. CLIP-I will be significantly reduced when image attention is removed. Simultaneously, the individual removal of image attention and structure attention impacts metrics such as FID and MUSIQ, which assess the quality of the generated images. Our method achieves a trade-off between realism and fidelity. Visually, it can be observed in Fig 9 that when removing image attention, the texture of the tire is not very clear. When removing structure attention, although the texture of the tire is clear, the structural details of the tire’s texture are incomplete. When removing image attention and structure attention, both the texture structure and clarity are compromised. Our method achieves better results in terms of structure and clarity thanks to both image attention and structure attention being employed.
We validate the directional restoration capability of our proposed method. Fig 10 provides an example. Given an LQ image, the generated restoration result using our model is shown in the second column, which is consistent with the text prompt generated by MLLM. Since the text output by MLLM is explicit, users can further modify the text details for personalized restoration. As shown in the third column, by changing the text prompt, the details in the restoration image can be effectively modified, which enhances the application value of the model.
We also discuss the generalizability of the proposed method. We provide more visual examples to show our method’s performance. As shown in Fig 11, our method shows great performance in various scenes, including old photos, faces, landscapes, outdoor and indoor photographs, and buildings. This proves that our method has wide application value.
Lastly, we analyze the complexity of the proposed method. The results are shown in Table VI. Compared with other methods, our method only needs one-stage training and without any other additional training. Since we introduced a new attention module, we inevitably introduced more training parameters and longer inference time. But the trainable parameters are still in the scale of hundred million and affordable.
V Conclusion And Limitations
@@
In this paper, we propose a novel multimodal-perception approach with semantic and structure priors for realistic image restoration based on the diffusion model. Regarding semantic information, we consider both explicit and implicit perspectives. We utilize the multimodal large language model to infer reliable semantic text representations from low-quality images. Additionally, we use the CLIP image encoder and a designed Refine Layer to generate the image embeddings, which collaborate with the text embedding to achieve a balance between fidelity and realism. For structure control, we employ the multi-level constrained Processor and ControlNet to extract structure conditions. Lastly, we integrate the control information and the diffusion model using the Prior-Guided Attention Module. Our method demonstrates state-of-the-art performance on multiple datasets both qualitatively and quantitatively.